### v1

# Beyond Face Swapping: A Diffusion-Based Digital Human Benchmark for Multimodal Deepfake Detection

Abstract.
In recent years, the rapid development of deepfake technology has given rise to an emerging and serious threat to public security: diffusion model-based digital human generation. Unlike traditional face manipulation methods, such models can generate highly realistic videos with consistency through multimodal control signals. Their flexibility and covertness pose severe challenges to existing detection strategies. To bridge this gap, we introduce DigiFakeAV [^1] https://hubeiwuhanliu.github.io/DigiFakeAV.github.io/ , the first large-scale multimodal digital human forgery dataset based on diffusion models. Employing five latest digital human generation methods (Sonic, Hallo, etc.) and voice cloning method, we systematically produce a dataset comprising 60,000 videos (8.4 million frames), covering multiple nationalities, skin tones, genders, and real-world scenarios, significantly enhancing data diversity and realism. User studies show that the confusion rate between forged and real videos reaches 68%, and existing state-of-the-art (SOTA) detection models exhibit large drops in AUC values on DigiFakeAV, highlighting the challenge of the dataset. To address this problem, we further propose DigiShield, a detection baseline based on spatiotemporal and cross-modal fusion. By jointly modeling the 3D spatiotemporal features of videos and the semantic-acoustic features of audio, DigiShield achieves SOTA performance on both the DigiFakeAV and DF-TIMIT datasets. Experiments show that this method effectively identifies covert artifacts through fine-grained analysis of the temporal evolution of facial features in synthetic videos.
Deepfake Detection, Multimodal Deepfake, Diffusion Model, Digital Human, Feature Fusion
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Computer vision
Figure 1. DigiFakeAV: A Multimodal Deepfake Dataset Based on Digital Human Technology.
[…]
DigiFakeAV.
1. Introduction
The rapid advancement of deepfake technology has posed unprecedented threats to digital security, giving rise to malicious applications such as political disinformation (Westerlund2019, ), synthetic pornographic content (Mirsky2021, ), and sophisticated financial fraud (yahoo2024hongkongscam, ). While existing detection benchmarks have driven progress, they remain constrained by outdated synthesis paradigms. DFDC (dolhansky2020deepfakedetectionchallengedfdc, ) proposes that deepfake detection datasets can be categorized into three generations. First-generation datasets (e.g., FF++ (rössler2019faceforensicslearningdetectmanipulated, )) suffer from limited scale and quality. The second generation (e.g., Celeb-DF (li2020celebdflargescalechallengingdataset, )) and third generation datasets (e.g., DFDC and FakeAVCeleb (khalid2022fakeavcelebnovelaudiovideomultimodal, )) further enhanced the quantity, authenticity, and diversity of data. However, third-generation datasets primarily focus on face-swapping methods (e.g., FSGAN, Faceswap) (kaur2024hindiaudiovideodeepfakehavdfhindi, ), leading to homogenized artifacts, audiovisual mismatches, and poor generalization to emerging threats.
Crucially, as diffusion models revolutionize synthetic media (chen2024echomimiclifelikeaudiodrivenportrait, ; lin2025omnihuman1rethinkingscalinguponestage, ; xu2024hallohierarchicalaudiodrivenvisual, ; cui2024hallo2longdurationhighresolutionaudiodriven, ; wang2024vexpressconditionaldropoutprogressive, ; ji2024sonicshiftingfocusglobal, ; tian2024emoemoteportraitalive, ), existing datasets fail to capture the multimodal realism and temporal coherence of modern digital human forgeries. As shown in Tab. 1, this has caused detection systems to approach performance saturation on current benchmarks while remaining fragile in real-world scenario (croitoru2024deepfakemediagenerationdetection, ). There is an urgent need for more challenging and representative datasets and benchmarks to address emerging threats.
To bridge this gap, we introduce DigiFakeAV, the first large-scale multimodal benchmark for detecting diffusion-based digital human forgeries. Our dataset advances the field through three key innovations:
Figure 2. Concrete examples demonstrating the diversity of DigiFakeAV, including multiple nationalities, diverse skin tones, various genders, and a range of scenarios.
$\bullet$ Diffusion-Driven Synthesis: Unlike previous benchmarks relying on GANs or face-swapping, DigiFakeAV employs five state-of-the-art diffusion models ( Sonic (ji2024sonicshiftingfocusglobal, ), Hallo (xu2024hallohierarchicalaudiodrivenvisual, ), etc.) to generate 60,000 high-resolution videos (8.4M frames). These models leverage multimodal signals (such as reference images and audio inputs) to create forgeries, eliminating source-target appearance constraints while achieving photorealistic details, such as subpixel skin textures.
$\bullet$ Multimodal Consistency: We ensure strict alignment between synthesized lip movements, facial expressions, and head poses with audio prosody and semantic content through speech-driven audiovisual synchronization. Compared with the baseline based on Wav2Lip, lip synchronization accuracy is significantly improved.
$\bullet$ Scene-Aware Diversity: DigiFakeAV covers various real-world scenarios (news broadcasts, social media vlogs, multilingual interviews) with balanced demographic representation: 57% male/43% female, as shown in Fig. 2. We also increase the proportion of Asians and Africans, which is in stark contrast to existing datasets, reducing racial bias while enhancing the robustness of cross-cultural detection.
Experimental validation reveals alarming vulnerabilities in current detectors: compared to face-swapping benchmarks, state-of-the-art models (e.g., SFIConv (10286083, )) exhibit a 43.5% drop in AUC on DigiFakeAV. Human evaluators misclassified 68% of synthetic videos as real, nearly triple the error rate of DFDC. These results underscore the urgency of developing detectors capable of identifying next-generation forgeries.
To advance research, we further propose DigiShield, a spatiotemporal audio-visual fusion framework that identifies synthetic artifacts through cross-modal inconsistency learning. By establishing the first effective baseline (AUC=80.1%), our work not only exposes critical gaps in detection research but also provides actionable insights for combating diffusion-based threats. DigiFakeAV represents a paradigm shift in deepfake benchmarking, prioritizing real-world relevance and adaptability to rapidly evolving generative technologies.
In summary, we conclude our work as follows:
$\bullet$ We introduce DigiFakeAV, the first large-scale multimodal benchmark for detecting diffusion-based digital human forgeries, comprising 60,000 high-resolution videos (8.4M frames) generated by five state-of-the-art diffusion models.
$\bullet$ Comprehensive experiments demonstrate that state-of-the-art detectors suffer a large AUC drop on DigiFakeAV, with human evaluators misclassifying 68% of synthetic videos as real, highlighting the urgency of advanced detection methods.
$\bullet$ We propose DigiShield, a spatiotemporal audiovisual fusion framework that identifies synthetic artifacts via cross-modal inconsistency learning, establishing an effective baseline for detecting next-level deepfakes.
Table 1. Performance (AUC) of state-of-the-art detection methods on existing deepfake video datasets. Results indicate that current benchmarks are no longer sufficiently challenging, motivating the creation of DigiFakeAV.
Generation Dataset Method AUC
1st UADFV (2018) Cross Efficient ViT(Coccomini_2022, ) 94.4
1st DF-TIMIT (2018) F3-Net(qian2020thinkingfrequencyfaceforgery, ) 99.8
1st FF++ (2019) TALL++(xu2024tallthumbnaillayoutdeepfake, ) 99.9
2nd Google DFD (2019) Cross Efficient ViT(Coccomini_2022, ) 93.0
2nd Celeb-DF (2020) App.+Beh.(9360904, ) 99.0
2nd DeeperForensics (2020) Style Latent Flow (choi2024exploitingstylelatentflows, ) 99.0
3rd DFDC (2020) FRADE(nie2024frade, ) 99.0
3rd KoDF (2021) AVFF(oorloff2024avffaudiovisualfeaturefusion, ) 95.5
3rd FakeAVCeleb (2021) FRADE(nie2024frade, ) 99.8
2. Related Work
2.1. Deepfake Detection Datasets
DFDC (dolhansky2020deepfakedetectionchallengedfdc, ) proposes that deepfake detection datasets can be categorized into three generations. The first-generation deepfake video datasets, such as DF-TIMIT, UADFV (yang2018exposingdeepfakesusing, ), and FaceForensics++, suffer from limitations in both scale and quantity—for example, UADFV contains only 98 videos. These early benchmark datasets laid the foundation for preliminary detection research, but are no longer applicable. The second-generation datasets, including Celeb-DF and DeeperForensics, have improved in scale and quality compared to the first generation, encompassing more forgery techniques and larger sample sizes. However, they still lack diversity in forgery types and multimodal synthetic content. Represented by DFDC and FakeAVCeleb, third-generation datasets include tens of thousands of samples and introduce audio synthesis technologies. DFDC contains synthetic but asynchronous audio, while FakeAVCeleb uses SV2TTS (jia2019transferlearningspeakerverification, ) and Wav2Lip for speech cloning and video reconstruction. Although speech synthesis technologies have been updated, other forgery techniques remain largely unchanged from second-generation datasets.
With the advancement of generative models, diffusion-based forgery has transcended facial regions and lip synchronization to involve multimodal digital human generation, achieving unprecedented levels of realism and audiovisual coherence. These technologies pose severe threats to information security, social trust, and media authenticity. To address this challenge, we created DigiFakeAV, a diffusion-based dataset featuring large-scale, high-quality, multimodal synthetic videos across diverse scenarios, providing a more comprehensive benchmark for robust deepfake detection.
2.2. Deepfake Video Generation
The progress of generative models has significantly advanced the evolution of deepfake videos. Early research in this field primarily relied on traditional GAN-based models for face swapping and reconstruction. For example, various multimodal forged video datasets employ Wav2Lip (Prajwal_2020, ), a method that utilizes a discriminator mechanism to improve the accuracy of audiovisual synchronization. Gao et al. (gao2023high, ) propose a face reenactment GAN focused on generating head movements synchronized with speech. However, GAN-based methods still face challenges such as unnatural head movements, facial distortions, and inaccurate identity retention. Additionally, training GANs requires a delicate balance between generators and discriminators, often leading to instability and sensitivity to data distributions.
In recent years, diffusion models have revolutionized the generation of videos based on portraits and digital humans. These methods take audio and reference images as input prompts, driving consistent and expressive animations through diffusion processes. For example, Animate Someone (hu2024animateanyoneconsistentcontrollable, ) can generate expressive videos with minimal control inputs. Based on this, EMO (tian2024emoemoteportraitalive, ) introduces a frame encoding module and improves temporal control to achieve coherent audio-driven animations. V-Express (wang2024vexpressconditionaldropoutprogressive, ) enhances audiovisual alignment through a hierarchical architecture, enabling precise synchronization between speech, lip movements, facial expressions, and head poses. Hallo (xu2024hallohierarchicalaudiodrivenvisual, ) proposes a hierarchical synthesis framework that effectively handles lip synchronization, expression, and pose alignment in portrait animation. In contrast, Sonic (ji2024sonicshiftingfocusglobal, ) adopts a global audio-aware design, achieving efficient long-video generation via time-shifted fusion without motion frames, while reducing computational overhead.
Synthetic videos generated by these diffusion-based methods are increasingly indistinguishable from real ones, marking the rise of next-level digital human-based forgery technologies and the obsolescence of current deepfake detection datasets. Therefore, there is an urgent need to establish new benchmarks based on diffusion models to drive corresponding advancements in detection and ensure their robustness against emerging forgery methods.
3. DigiFakeAV
In this section, we describe the pipeline of dataset construction, including data collection, data synthesis, and the composition of the final dataset, as shown in Fig. 3. In the core subsection of data synthesis, we discuss the specific generation methods.
Figure 3. Workflow Diagram for the Construction of DigiFakeAV.
[…]
Workflow Diagram
3.1. Data Collection
The real videos used in this study are sourced from the HDTF (zhang2021flow, ) and CelebV-HQ (zhu2022celebvhqlargescalevideofacial, ) datasets. We implement a dedicated video processing pipeline to split all MP4, AVI, and MKV files in the original directories into shorter clips, which are then cropped and re-encoded. Specifically, for each video file, we first calculate its total duration and split it into segments of 3–5 seconds each.
Subsequently, we use FFmpeg to process each clip, including resizing to 512×512 pixels, copying the audio stream without re-encoding, and setting the bitrate to 3000 kbps to balance processing speed and quality. Through the above steps, we obtain nearly 40,000 clean video clips from the HDTF and CelebV-HQ datasets for further analysis.
3.2. Data Synthesis
We produce the DigiFakeAV dataset using state-of-the-art deepfake and audio synthesis models. Specifically, we employ five recent portrait animation video generation methods and one advanced audio synthesis method. As summarized in Tab. 2, compared with previous deepfake generation methods, our choices exhibit significant differences and advantages.
Synthesis Pipeline. Both methods generate fake video and fake audio samples through three stages: (1) Generation conditioning: The video generation model is driven by real audio (from real video–real audio clips) or synthetic audio (e.g., TTS), paired with static portraits or video frames as visual inputs. (2) Diffusion sampling: Video frames are synthesized via iterative denoising guided by audio and facial priors. Clips are 5–10 seconds long, meeting the dataset’s resolution (512×512) and bitrate (3000 kbps) requirements. (3) Quality control: Samples are filtered using perceptual metrics (e.g., FID, Sync-C) to eliminate artifacts such as lip-sync errors or identity drift. Realistic noise and compression are added to simulate real-world distribution shifts.
Key Advantages. (1) Higher Realism: Diffusion models can generate fine-grained facial details (e.g., skin textures, eye movements) and maintain long-term temporal consistency, outperforming GAN-based methods. (2) Audiovisual Consistency: Sonic and Hallo achieve strong lip-sync accuracy and emotional consistency, challenging unimodal detection. (3) Scalability: Both frameworks support batch processing and domain-specific fine-tuning (e.g., celebrity voice dubbing), enabling rapid dataset expansion.
By integrating these methods, DigiFakeAV provides rich FV-FA samples that closely resemble real-world forgery scenarios, thus supporting the development of robust audiovisual deepfake detection systems.
3.3. Dataset Organization
Table 2. Quantitative Comparison of the DigiFakeAV Dataset with Previous Datasets
Dataset Year Real / Fake Total Modality Methods Real Audio Deepfake Audio UADFV (yang2018exposingdeepfakesusing, ) 2018 49 / 49 98 V 1 No No DF-TIMIT (korshunov2018deepfakesnewthreatface, ) 2018 320 / 640 960 V 2 Yes No FF++ (rössler2019faceforensicslearningdetectmanipulated, ) 2019 1,000 / 4,000 5,000 V 4 No No Celeb-DF (li2020celebdflargescalechallengingdataset, ) 2020 590 / 5,639 6,229 V 1 No No Google DFD (rössler2019faceforensicslearningdetectmanipulated, ) 2019 363 / 3,000 3,363 V 5 No No DeeperForensics (jiang2020deeperforensics10largescaledatasetrealworld, ) 2020 50,000 / 10,000 60,000 V 1 No No DFDC (dolhansky2020deepfakedetectionchallengedfdc, ) 2020 23,654 / 104,500 128,154 A/V 8 Yes Yes KoDF (Kwon_2021_ICCV, ) 2021 62,166 / 175,776 237,942 V 6 Yes No FakeAVCeleb (khalid2022fakeavcelebnovelaudiovideomultimodal, ) 2021 500 / 19,500 20,000 A/V 4 Yes Yes DeepSpeak (barrington2025deepspeakdataset, ) 2024 6,226 / 6,799 13,025 A/V 4 Yes Yes DigiFakeAV (ours) 2025 10,000 / 50,000 60,000 A/V 6 Yes Yes
Forged videos and audio can form four combinations: (1) real video–real audio (RV-RA), (2) real video–fake audio (RV-FA), (3) fake video–real audio (FV-RA), and (4) fake video–fake audio (FV-FA). We exclude RV-FA for two reasons: First, our focus is on video-based facial forgery detection, whereas RV-FA does not involve human facial manipulation. For example, DFDC (dolhansky2020deepfakedetectionchallengedfdc, ) labels RV-FA samples as fake, which is inherently unfair to vision-only detectors. Second, Audio-only manipulations without visual changes are often obvious and pose limited detection challenges. For example, FakeAVCeleb (khalid2022fakeavcelebnovelaudiovideomultimodal, ) contains audio-video synchronization delays.
We then introduce three combinations and the corresponding methods adopted in this study, with detailed explanations in subsequent use cases:
Real Video–Real Audio (RV-RA). From nearly 40,000 clean video clips, we select 10,000 real videos representing diverse ethnicities, genders, and ages. These videos are further processed to crop portrait regions, forming the basis of our real video dataset.
Fake Video–Real Audio (FV-RA) This category contains fake videos generated by synthesizing visual content conditioned on real audio. We first extract and convert audio from real videos into WAV format. RetinaFace is then used to detect and crop representative facial frames. Using five digital human generation methods: Sonic (ji2024sonicshiftingfocusglobal, ), Hallo1 (xu2024hallohierarchicalaudiodrivenvisual, ), Hallo2 (cui2024hallo2longdurationhighresolutionaudiodriven, ), Echomimic (chen2024echomimiclifelikeaudiodrivenportrait, ), and V-Express (wang2024vexpressconditionaldropoutprogressive, ), we generate a total of 25,000 forged videos. Modern attackers can acquire the speech and facial data of a target to conduct highly realistic impersonation attacks, surpassing the threats of traditional face-swapping methods.
Fake Video–Fake Audio (FV-FA). This category encompasses both forged audio and video. We employed a state-of-the-art voice cloning method, CosyVoice 2 (du2024cosyvoice2scalablestreaming, ). Specifically, we first use a large language model to generate processed text. By providing real audio-text pairs, the CosyVoice 2 model synthesizes fake audio mimicking the target speaker’s voice characteristics. We then generate 25,000 forged videos based on this fake audio using algorithms like Sonic, Hallo1, Hallo2, and Echomimic. This subset helps researchers study advanced identity deception scenarios.
3.4. Dataset Statistics
We note that existing datasets (e.g., HDTF and CelebV-HQ) may exhibit gender and ethnic biases. To address this, we prioritize diversity across four dimensions during video selection(Fig. 2): (1) Gender balance: We achieve a 57:43 male-female ratio to reflect real-world demographics and avoid excessive gender skew. (2) Ethnic diversity: The dataset includes White, African American, and Asian individuals, ensuring broad racial representation and enhancing model generalization. (3) Nationality and cultural background: Subjects from diverse countries and cultural contexts are selected to reduce regional bias and enhance global representativeness. (4) Scenario diversity: Video content simulates real-world scenarios, including social interactions, news broadcasts, and daily conversations, to improve applicability in practical contexts.
3.5. User Study
To evaluate the quality of the DigiFakeAV dataset, we recruited 100 participants with expertise in computer vision research, as we deemed their experience and domain knowledge essential for effectively assessing forged videos. The user study covers DigiFakeAV and three prior datasets—FF-DF, Celeb-DF, and DFDC—representing the first, second, and third-generation deepfake video benchmarks, respectively.
We select 50 video clips from each dataset and provide participants with a platform to evaluate each video’s authenticity. Participants are asked to vote on whether each clip was real or fake. The results of the user study, summarized in Tab.  3, clearly demonstrate that the proposed dataset was rated as highly challenging, highlighting the realism and detection difficulty of synthetic videos in DigiFakeAV.
Table 3. User Ratings of Fake Video Authenticity Across Datasets.
Dataset FF-DF Celeb-DF DFDC DigiFakeAV (ours) Fake / Real (%) 12% 46% 32% 68%
4. DigiShield
We propose DigiShield, a multimodal audio-visual spatiotemporal fusion detector. Our approach focuses on identifying subtle variations in facial features over time in synthetic videos. To capture such fine-grained temporal inconsistencies, we employ 3D convolution to model spatial and temporal patterns jointly. The integration of audio further enhances robustness by providing synchronization cues that improve forgery localization.
4.1. Model Framework
We find that single-modality analysis is insufficient for detecting digital human forgery artifacts. Thus, it is crucial to jointly model spatial and temporal inconsistencies without over-relying on any single cue. Building on prior 3D CNN and two-stream approaches (wang2023altfreezinggeneralvideoface, ; masi2020twobranchrecurrentnetworkisolating, ), we further exploit audio-visual cues through multimodal fusion. Our method includes (1) a spatiotemporal two-stream pipeline, (2) audio-visual fusion, and (3) a classification layer (see Fig.4). The model is trained using both contrastive and cross-entropy losses to capture audio-visual inconsistencies indicative of forgery.
Spatiotemporal two-stream pipeline. Formally, given a set of video clips $x_{\raisebox{-0.90417pt}{\scriptsize in}}^{\raisebox{0.60275pt}{\scriptsize v}}\in\mathbb{R}^{T\times C\times H_{v}\times W_{v}}$ and audio signals $x_{\raisebox{-0.90417pt}{\scriptsize in}}^{\raisebox{0.60275pt}{\scriptsize a}}\in\mathbb{R}^{C\times H_{a}\times W_{a}}$ (where T denotes the number of frames in the same video, C represents the number of channels, and H × W indicates the spatial dimensions), we first employ a spatiotemporal two-stream pipeline based on 2D and 3D residual convolutions to capture forgery traces. The spatiotemporal representations of the video and audio can then be computed as follows:
(1) $f_{v}=\mathcal{F}_{v}\biggl{(}x_{\raisebox{-0.90417pt}{\scriptsize in}}^{\raisebox{0.60275pt}{\scriptsize v}}\biggr{)}\in\mathbb{R}^{T^{\prime}\times C^{\prime}\times H^{\prime}_{v}\times W^{\prime}_{v}},f_{a}=\mathcal{F}_{a}\biggl{(}x_{\raisebox{-0.90417pt}{\scriptsize in}}^{\raisebox{0.60275pt}{\scriptsize a}}\biggr{)}\in\mathbb{R}^{C^{\prime}\times H^{\prime}_{a}\times W^{\prime}_{a}}$
The variables $T^{\prime}$,$C^{\prime}$,$H^{\prime}$, and $W^{\prime}$denote the temporal, channel, height, and width dimensions, respectively, after processing through the convolution.
Multimodal spatiotemporal fusion. The spatiotemporal features of video and audio, denoted as $\mathbf{f}_{v}\in\mathbb{R}^{N_{v}\times d}$ and $\mathbf{f}_{a}\in\mathbb{R}^{N_{a}\times d}$ respectively, are extracted through the video encoder $\mathcal{F}_{v}$ and audio encoder $\mathcal{F}_{a}$, and subsequently flattened along their spatiotemporal dimensions. Here, $N_{v}$ and $N_{a}$ denote the number of spatiotemporal tokens for video and audio, respectively, and $d$ represents the feature dimension.The flattened features are denoted as $\widetilde{\mathbf{f}}_{v}\in\mathbb{R}^{(N_{v}d)}$ and $\widetilde{\mathbf{f}}_{a}\in\mathbb{R}^{(N_{a}d)}$, which serve as input tensors to the multi-head attention mechanism.
We then apply cross-attention and self-attention modules to enhance multimodal interaction and feature fusion as follows:
(2) $\Psi$ $=\operatorname{CrossAtt}(\widetilde{f}_{v},\widetilde{f}_{a})=\operatorname{softmax}\left(\frac{\widetilde{f}_{v}W_{Q}(\widetilde{f}_{a}W_{K})^{\top}}{\sqrt{d_{k}}}\right)\widetilde{f}_{a}W_{V},$
$\Phi$ $=\operatorname{SelfAtt}(\Psi)=\operatorname{softmax}\left(\frac{\Psi W_{Q}^{\prime}(\Psi W_{K}^{\prime})^{\top}}{\sqrt{d_{k}}}\right)\Psi W_{V}^{\prime}.$
where $W_{Q},W_{K},W_{V}\in\mathbb{R}^{d\times d_{k}}$ are the learned projection matrices for queries, keys, and values in the cross-attention module; $W_{Q}^{\prime},W_{K}^{\prime},W_{V}^{\prime}\in\mathbb{R}^{d\times d_{k}}$ are the corresponding projection matrices in the self-attention module; $d_{k}$ is the dimension of the key (and query) vectors used for scaling the dot-product attention.
Final decision layer. The fused feature obtained is mapped to a binary classification probability space to make the final decision on the input video.
Figure 4. Overview of the DigiShield model: it detects forgery traces by integrating spatiotemporal information from both video and audio modalities.
4.2. Loss Function
To effectively fuse visual and audio modal features, we adopt a contrastive loss defined as:
(3) $\small\mathcal{L}_{con}=\frac{1}{N}\sum_{i=1}^{N}\left[y_{i}\,D(\widetilde{\mathbf{h}}_{v}^{(i)},\widetilde{\mathbf{h}}_{a}^{(i)})^{2}+(1-y_{i})\,\max\big{(}0,m-D(\widetilde{\mathbf{h}}_{v}^{(i)},\widetilde{\mathbf{h}}_{a}^{(i)})\big{)}^{2}\right]$
where $N$ is the batch size, $\widetilde{\mathbf{h}}_{v}^{(i)}$ and $\widetilde{\mathbf{h}}_{a}^{(i)}$ are the flattened visual and audio features of the $i$-th sample, $y_{i}\in\{0,1\}$ denotes ground-truth labels (1 for genuine, 0 for fake), $D(\cdot,\cdot)$ is the Euclidean distance, i.e., $D(\mathbf{x},\mathbf{y})=\|\mathbf{x}-\mathbf{y}\|_{2}$, and $m>0$ is a margin parameter.
The classification is guided by cross-entropy loss:
(4) $\mathcal{L}_{ce}=-\frac{1}{N}\sum_{i=1}^{N}\left[y_{i}\log\hat{y}_{i}+(1-y_{i})\log(1-\hat{y}_{i})\right]$
where $\hat{y}_{i}\in[0,1]$ is the predicted probability of the $i$-th sample being genuine.
The overall loss is a weighted sum:
(5) $\mathcal{L}=\mathcal{L}_{con}+\mathcal{L}_{ce}$
Here, ${L}_{con}$ and $\mathcal{L}_{ce}$ denote the contrastive loss and the cross-entropy loss, respectively.
5. Experiment
5.1. Implementation Details
Preprocessing. During the preprocessing stage, we perform specialized processing on video and audio data. For the video modality, we use the RetinaFace face detection model to locate facial regions and perform corresponding cropping. For the audio modality, we first extract audio from videos, resample it to 16 kHz, and store it in WAV format. We then compute the Mel-frequency cepstral coefficients (MFCC) for each audio frame and convert them into a three-channel image representation as the audio input for the multimodal model.
Evaluation Protocol. For DigiShield, we adopt 3D-ResNet-50 as the default backbone. To ensure a fair and comprehensive evaluation of our proposed dataset, all experiments are conducted on data splits that follow an 8:1:1 ratio for training, validation, and testing, respectively, with no identity overlap between training and testing subsets. During training, we uniformly sample 20 frames from DF-TIMIT and 30 frames from DigiFakeAV videos, and apply data augmentation including color jittering and random cropping. To address the class imbalance in DigiFakeAV, we further apply heuristic upsampling to real samples during training, ensuring a balanced distribution between real and fake data. The AUC was employed as the primary evaluation metric. All experiments were conducted under consistent data preprocessing and training conditions to guarantee the stability and reliability of the findings.
Table 4. AUC scores (%) of various methods on compared datasets.
Method DF-TIMIT FF-DF DFDC Celeb-DF FakeAVCeleb DigiFakeAV LQ HQ (ours) Meso4 87.8 68.4 84.7 75.3 54.8 60.9 50.1 MesoInception4 80.4 62.7 83.0 73.2 53.6 61.7 63.8 Xception-c23 95.9 94.4 99.7 72.2 65.3 72.5 66.1 Capsule 78.4 74.4 96.6 53.3 57.5 70.9 65.3 HeadPose 55.1 53.2 47.3 55.9 54.6 49.0 51.7 F3-Net 99.8 99.4 93.7 95.1 86.7 91.3 66.4 Cross Efficient ViT 50.4 55.8 99.1 95.1 86.7 80.5 52.2 SSVF - - - - - 94.5 51.0 SFIConv 100.0 100.0 95.9 96.7 95.8 93.0 71.2
5.2. Baselines Comparison
To conduct a thorough evaluation of the DigiFakeAV dataset, we systematically select nine representative deepfake detection algorithms based on the criteria of representativeness and accessibility. These methods encompass six distinct categories: Basic (Naive), Spatial, Temporal, Frequency, Hybrid Domain, and Multi-modal approaches. The chosen models comprise Meso4 (Afchar_2018, ), MesoInception4 (Afchar_2018, ), Xception-c23 (rössler2019faceforensicslearningdetectmanipulated, ), Capsule (nguyen2019usecapsulenetworkdetect, ), HeadPose (yang2018exposingdeepfakesusing, ), F3-Net (qian2020thinkingfrequencyfaceforgery, ), Cross Efficient ViT (Coccomini_2022, ), SFIConv (10286083, ), and SSVF (feng2023selfsupervisedvideoforensicsaudiovisual, ).
We follow standard benchmarking protocols, with the precision setting for baseline experiments rounding all classification scores to three decimal places ( $\ 10^{-3}$). As shown in Tab. 4, most evaluated methods exhibit significant performance degradation on DigiFakeAV compared to their results on other public datasets. For example, the AUC of SFIConv drops from 100.0% (DF-TIMIT HQ) to 71.2%, and SSVF declines from 94.5% (FakeAVCeleb) to 51.0%. These results highlight the effectiveness of digital human generation technologies in evading detection and expose the vulnerability of current models to advanced forgery strategies. A comprehensive description of the benchmark methods employed and further experimental evaluations are provided in the Supplementary Materials for reference.
5.3. DigiShield Comparison
As shown in Tab.  5, our model achieves an AUC of 80.1 on the more challenging DigiFakeAV dataset, outperforming existing methods and indicating that the digital human forgeries we propose require richer multimodal cues for effective detection. Our approach also achieves state-of-the-art (SOTA) performance on the DF-TIMIT dataset with an AUC score of 1.0, demonstrating its strong generalization ability across data of varying qualities.
Table 5. Comparison of AUC scores for DigiShield on DF-TIMIT and DigiFakeAV.
Method DF-TIMIT-LQ DF-TIMIT-HQ DigiFakeAV
MesoInception4 80.4 62.7 63.8
Capsule 78.4 74.4 65.3
Xception-c23 95.9 94.4 66.1
F3-Net 99.8 99.4 66.4
SFIConv 100.0 100.0 71.2
DigiShield 100.0 100.0 80.1
Table 6. Ablation study of AVTSF in DigiFakeAV.
$\mathcal{L}_{ce}$ $\mathcal{L}_{con}$ CrossAtt SelfAtt AUC
73.6
77.4
80.1
5.4. Ablation Study
Tab. 6 validates the effectiveness of our audio-visual spatiotemporal fusion strategy. Using only cross-entropy loss ($L_{ce}$) achieves an AUC of 73.6, while adding contrastive loss ($L_{con}$) and Cross-attention (CrossAtt) improves it to 77.4. Self-attention (SelfAtt) further enhances temporal coherence, achieving an AUC of 80.1.
These results highlight the importance of both inter-modal and intra-modal attention mechanisms in detecting subtle digital human artifacts and diffusion-based forgery. For more detailed information, please refer to the Supplementary Materials.
6. Conclusion
DigiFakeAV is a novel audio-visual multimodal dataset based on advanced digital human forgery techniques. As the first dataset emphasizing diffusion-based synthesis, it introduces unprecedented challenges for deepfake detection. Its difficulty is evaluated using nine detection methods, revealing significant performance drops compared to previous benchmarks. Additionally, DigiShield, an audio-visual fusion baseline, is proposed to support future research. DigiFakeAV is expected to foster the development of more robust deepfake detectors and inspire new directions in detection research.