### v2

# Beyond Face Swapping: A Diffusion-Based Digital Human Benchmark for Multimodal Deepfake Detection

Abstract.
In recent years, the rapid development of deepfake technology has given rise to an emerging and serious threat to public security: diffusion-based digital human generation. Unlike traditional face manipulation methods, such models can generate highly realistic videos with consistency through multimodal control signals. Their flexibility and covertness pose severe challenges to existing detection strategies. To bridge this gap, we introduce DigiFakeAV [^1] https://hubeiwuhanliu.github.io/DigiFakeAV.github.io/, the new large-scale multimodal digital human forgery dataset based on diffusion models. Employing five of the latest digital human generation methods and the voice cloning methods, we systematically produce a dataset comprising 60,000 videos (8.4 million frames), covering multiple nationalities, skin tones, genders, and real-world scenarios, significantly enhancing data diversity and realism. User studies demonstrate that participants misclassify forged videos as real in 68% of tests, and existing detection models exhibit a large drop in performance on DigiFakeAV, highlighting the challenge of the dataset. To address this problem, we propose DigiShield, an effective detection baseline based on spatiotemporal and cross-modal fusion. By jointly modeling the 3D spatiotemporal features of videos and the semantic-acoustic features of audio, DigiShield achieves state-of-the-art (SOTA) performance on the DigiFakeAV and shows strong generalization on other datasets.
Deepfake Detection, Multimodal Deepfake, Diffusion Model, Digital Human, Feature Fusion
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Conference’33; October 2025; Dublin, Irelandbooktitle: Conference’33, 2025, Dublin, Irelandisbn: 978-1-4503-XXXX-X/2018/06ccs: Security and privacy Human and societal aspects of security and privacy
Figure 1. DigiFakeAV: A Multimodal Deepfake Dataset Based on Digital Human Technology.
[…]
DigiFakeAV.
1. Introduction
The rapid advancement of deepfake technology has posed unprecedented threats to digital security, giving rise to malicious applications such as political disinformation (Westerlund2019, ), synthetic pornographic content (Mirsky2021, ), and sophisticated financial fraud (yahoo2024hongkongscam, ). While existing detection benchmarks have driven progress, they remain constrained by outdated synthesis paradigms. The existing deepfake detection datasets can be roughly categorized into three generations (dolhansky2020deepfakedetectionchallengedfdc, ). First-generation datasets (e.g., FF++ (rössler2019faceforensicslearningdetectmanipulated, )) suffer from limited scale and quality. The second generation (e.g., Celeb-DF (li2020celebdflargescalechallengingdataset, )) and third generation datasets (e.g., DFDC (dolhansky2020deepfakedetectionchallengedfdc, )) further enhanced the quantity, authenticity, and diversity of data. However, the latest third-generation datasets primarily focus on face-swapping methods (e.g., FSGAN, Faceswap) (kaur2024hindiaudiovideodeepfakehavdfhindi, ), leading to homogenized artifacts, audiovisual mismatches, and poor generalization to emerging threats.
Crucially, as diffusion models revolutionize synthetic media (chen2024echomimiclifelikeaudiodrivenportrait, ; lin2025omnihuman1rethinkingscalinguponestage, ; xu2024hallohierarchicalaudiodrivenvisual, ; cui2024hallo2longdurationhighresolutionaudiodriven, ; wang2024vexpressconditionaldropoutprogressive, ; ji2024sonicshiftingfocusglobal, ; tian2024emoemoteportraitalive, ), existing datasets fail to capture the multimodal realism and temporal coherence of modern digital human forgeries. As shown in Tab. 1, this has caused detection systems to approach performance saturation on current benchmarks while remaining fragile in real-world scenarios (croitoru2024deepfakemediagenerationdetection, ). There is an urgent need for more challenging and representative datasets and benchmarks to address emerging threats.
To bridge this gap, we introduce DigiFakeAV, the new large-scale multimodal benchmark for detecting diffusion-based digital human forgeries. Our dataset advances the field through three key innovations:
Figure 2. Examples demonstrating the diversity of DigiFakeAV, including multiple nationalities, diverse skin tones, various genders, and a range of scenarios.
[…]
Examples demonstrating the diversity of DigiFakeAV
$\bullet$ Diffusion-Driven Synthesis: unlike previous benchmarks relying on GANs or face-swapping, DigiFakeAV employs five state-of-the-art diffusion models ( Sonic (ji2024sonicshiftingfocusglobal, ), Hallo (xu2024hallohierarchicalaudiodrivenvisual, ), etc.) to generate 60,000 high-resolution videos (8.4 million frames). These models leverage multimodal signals (such as reference images and audio inputs) to create forgeries, eliminating source-target appearance constraints while achieving photorealistic details, such as subpixel skin textures.
$\bullet$ Multimodal Consistency: we ensure strict alignment between synthesized lip movements, facial expressions, and head poses with audio prosody and semantic content through speech-driven audiovisual synchronization. Lip synchronization is significantly improved compared with the previous datasets based on Wav2Lip (Prajwal_2020, ).
$\bullet$ Scene-Aware Diversity: DigiFakeAV covers various real-world scenarios (news broadcasts, social media vlogs, multilingual interviews) with balanced demographic representation: 57% male/43% female, as shown in Fig. 2. We also increase the proportion of Asians and Africans, reducing bias while enhancing the robustness of cross-cultural detection.
Experiments reveal alarming vulnerabilities in current detectors: compared to face-swapping benchmarks, SOTA models experience a dramatic drop in Area Under the Curve (AUC) on DigiFakeAV. For example, SSVF (feng2023selfsupervisedvideoforensicsaudiovisual, ) exhibits a 43.5% decrease. Human evaluators misclassified 68% of synthetic videos as real, nearly triple the error rate of DFDC (dolhansky2020deepfakedetectionchallengedfdc, ). This suggests that we must design detectors that can identify next-generation forgeries.
To address the challenge of diffusion-based digital human forgeries, we propose DigiShield, a new multimodal detection framework. The framework explicitly captures dynamic inconsistencies between video and audio through cross-modal attention mechanisms and spatiotemporal feature extraction networks, enabling effective identification of subtle forgery traces generated by diffusion models. Experiments show that DigiShield outperforms existing methods on the DigiFakeAV dataset, validating its effectiveness in next-generation deepfake detection and providing scalable framework design insights for future research.
In summary, we conclude our work as follows:
$\bullet$ We introduce DigiFakeAV, the new large-scale multimodal benchmark for detecting diffusion-based digital human forgeries, comprising 60,000 high-resolution videos (8.4 million frames) generated by SOTA diffusion models.
$\bullet$ Comprehensive experiments demonstrate that SOTA detectors suffer a large performance drop on DigiFakeAV, with human evaluators misclassifying 68% of synthetic videos as real, highlighting the urgency of advanced detection methods.
$\bullet$ We propose DigiShield, a spatiotemporal audiovisual fusion framework that identifies synthetic artifacts via cross-modal inconsistency learning, establishing a new baseline for detecting next-level deepfakes.
Table 1. Performance (AUC) of SOTA detection methods on existing deepfake video datasets. Results indicate that current benchmarks are constrained by outdated synthesis paradigms.
Generation Dataset Method AUC
1st UADFV (2018) Cross Efficient ViT (Coccomini_2022, ) 94.4
1st DF-TIMIT (2018) F3-Net (qian2020thinkingfrequencyfaceforgery, ) 99.8
1st FF++ (2019) TALL++ (xu2024tallthumbnaillayoutdeepfake, ) 99.9
2nd Google DFD (2019) Cross Efficient ViT (Coccomini_2022, ) 93.0
2nd Celeb-DF (2020) App.+Beh. (9360904, ) 99.0
2nd DeeperForensics (2020) Style Latent Flow (choi2024exploitingstylelatentflows, ) 99.0
3rd DFDC (2020) FRADE (nie2024frade, ) 99.0
3rd KoDF (2021) AVFF (oorloff2024avffaudiovisualfeaturefusion, ) 95.5
3rd FakeAVCeleb (2021) FRADE (nie2024frade, ) 99.8
Table 2. Quantitative comparison of the DigiFakeAV dataset with previous datasets
Dataset Year Real / Fake Total Modality Methods Real Audio Deepfake Audio UADFV (yang2018exposingdeepfakesusing, ) 2018 49 / 49 98 V 1 No No DF-TIMIT (korshunov2018deepfakesnewthreatface, ) 2018 320 / 640 960 V 2 Yes No FF++ (rössler2019faceforensicslearningdetectmanipulated, ) 2019 1,000 / 4,000 5,000 V 4 No No Celeb-DF (li2020celebdflargescalechallengingdataset, ) 2020 590 / 5,639 6,229 V 1 No No Google DFD (rössler2019faceforensicslearningdetectmanipulated, ) 2019 363 / 3,000 3,363 V 5 No No DeeperForensics (jiang2020deeperforensics10largescaledatasetrealworld, ) 2020 50,000 / 10,000 60,000 V 1 No No DFDC (dolhansky2020deepfakedetectionchallengedfdc, ) 2020 23,654 /104,500 128,154 A/V 8 Yes Yes KoDF (Kwon_2021_ICCV, ) 2021 62,166 /175,776 237,942 V 6 Yes No FakeAVCeleb (khalid2022fakeavcelebnovelaudiovideomultimodal, ) 2021 500 /19,500 20,000 A/V 4 Yes Yes DeepSpeak (barrington2025deepspeakdataset, ) 2024 6,226 / 6,799 13,025 A/V 4 Yes Yes DigiFakeAV (ours) 2025 10,000 /50,000 60,000 A/V 6 Yes Yes
2. Related Work
2.1. Deepfake Detection Datasets
DFDC (dolhansky2020deepfakedetectionchallengedfdc, ) proposes that deepfake detection datasets can be roughly categorized into three generations. The first-generation deepfake video datasets, such as DF-TIMIT (korshunov2018deepfakesnewthreatface, ), UADFV (yang2018exposingdeepfakesusing, ), and FF++  (rössler2019faceforensicslearningdetectmanipulated, ), suffer from limitations in both scale and quantity. For example, UADFV contains only 98 videos. These early benchmark datasets laid the foundation for initial detection research, but are no longer applicable. The second-generation datasets, including Celeb-DF (li2020celebdflargescalechallengingdataset, ) and DeeperForensics (jiang2020deeperforensics10largescaledatasetrealworld, ), have improved in scale and quality compared to the first generation, encompassing more forgery techniques and larger sample sizes. However, they still lack diversity in forgery types and multimodal synthetic content. Represented by DFDC (dolhansky2020deepfakedetectionchallengedfdc, ) and FakeAVCeleb (khalid2022fakeavcelebnovelaudiovideomultimodal, ), third-generation datasets include tens of thousands of samples and introduce audio synthesis technologies. DFDC contains synthetic but asynchronous audio, while FakeAVCeleb uses SV2TTS (jia2019transferlearningspeakerverification, ) and Wav2Lip (Prajwal_2020, ) for speech cloning and video reconstruction. Except for speech synthesis technologies, other forgery techniques remain unchanged from the second generation.
With advancements in generative models, diffusion-based forgery (wei2024aniportraitaudiodrivensynthesisphotorealistic, ) now enables highly realistic and coherent multimodal digital human generation beyond facial and lip synthesis. To address the detection challenges this poses, we introduce DigiFakeAV, a high-quality, large-scale diffusion-based dataset of multimodal synthetic videos spanning diverse scenarios, offering a comprehensive benchmark for robust deepfake detection.
2.2. Deepfake Video Generation
The progress of generative models has significantly advanced the evolution of deepfake videos. Early research primarily relies on traditional GAN-based models for face swapping and reconstruction. For example, various multimodal forged video datasets employ Wav2Lip (Prajwal_2020, ), utilizing a discriminator mechanism to improve audiovisual synchronization. Gao et al. (gao2023high, ) propose a face reenactment GAN focused on generating head movements synchronized with speech. However, GAN-based methods still face challenges such as unnatural head movements, facial distortions, and inaccurate identity retention. Additionally, training GANs often leads to instability and sensitivity to data distributions.
In recent years, diffusion models have revolutionized the generation of digital human videos. These methods take audio and reference images as input prompts, driving consistent and expressive animations through diffusion processes. For example, Animate Someone  (hu2024animateanyoneconsistentcontrollable, ) can generate expressive videos with minimal control inputs. Based on this, EMO (tian2024emoemoteportraitalive, ) introduces a frame encoding module and improves temporal control to achieve coherent audio-driven animations. V-Express (wang2024vexpressconditionaldropoutprogressive, ) enhances audiovisual alignment through a hierarchical architecture. Hallo (xu2024hallohierarchicalaudiodrivenvisual, ) proposes a hierarchical synthesis framework that effectively handles lip synchronization. Moreover, Sonic (ji2024sonicshiftingfocusglobal, ) adopts a global audio-aware design, achieving efficient long-video generation via time-shifted fusion without motion frames.
Synthetic videos generated by these diffusion-based methods are indistinguishable from real ones, marking the rise of next-level forgery technologies and the obsolescence of current deepfake detection datasets. Therefore, there is an urgent need to establish new benchmarks based on diffusion models to drive corresponding advancements in detection and ensure their robustness against emerging forgery methods.
3. DigiFakeAV: A Next Generation Deepfake Benchmark
In this section, we describe the pipeline of dataset construction, including data collection, data synthesis, and the composition of the final dataset, as shown in Fig. 3. In the core subsection of data synthesis, we discuss the specific generation methods [^2] For further details of our dataset, please refer to the Supplementary Materials (Appendix A), at https://github.com/DigiFakeAV/DigiFakeAV-v1/blob/main/supplementary/Supplementary.pdf .
Figure 3. Workflow diagram for the construction of DigiFakeAV.
[…]
Workflow Diagram
3.1. Data Collection
The real videos of our dataset are sourced from the HDTF (zhang2021flow, ) and CelebV-HQ (zhu2022celebvhqlargescalevideofacial, ) datasets. We implement a dedicated video processing pipeline to split all MP4, AVI, and MKV files in the original directories into shorter clips, which are then cropped and re-encoded. Specifically, for each file, we first calculate its total duration and split it into segments of 3–5 seconds each. Subsequently, we use FFmpeg to process each clip, including resizing to 512×512 pixels, copying the audio stream without re-encoding, and setting the bitrate to 3000 kbps to balance processing speed and quality.
Through the above steps, we obtain nearly 40,000 clean video clips from the HDTF and CelebV-HQ for further analysis.
3.2. Data Synthesis
We produce the DigiFakeAV dataset using state-of-the-art (SOTA) deepfake and audio synthesis models. Specifically, we employ five recent diffusion-based digital human video generation methods (Sonic (ji2024sonicshiftingfocusglobal, ), Hallo (xu2024hallohierarchicalaudiodrivenvisual, ), Hallo2 (cui2024hallo2longdurationhighresolutionaudiodriven, ), EchoMimic (chen2024echomimiclifelikeaudiodrivenportrait, ), and V-Express (wang2024vexpressconditionaldropoutprogressive, )) and one advanced audio synthesis method (CosyVoice 2 (du2024cosyvoice2scalablestreaming, )). As summarized in Tab. 2, compared with previous deepfake methods, our choices exhibit significant advantages.
Synthesis Pipeline. Both methods generate fake video and fake audio samples through three stages: (1) Generation conditioning: the video generation model is driven by real audio (from real video–real audio clips) or synthetic audio (e.g., TTS), paired with static portraits or video frames as visual inputs. (2) Diffusion sampling: video frames are synthesized via iterative denoising guided by audio and facial priors. Clip lengths range from 5 to 10 seconds, meeting the dataset’s resolution (512×512) and bitrate (3000 kbps) requirements. (3) Quality control: samples are filtered using perceptual metrics (e.g., FID, Sync-C) to eliminate artifacts such as lip-sync errors or identity drift [^3] Generation quality Standards are in Appendix A.3. Realistic noise and compression are added to simulate real-world distribution shifts.
Key Advantages. (1) Higher Realism: diffusion models can generate fine-grained facial details (skin textures, eye movements) and maintain long-term temporal consistency, outperforming GAN-based methods. (2) Audiovisual Consistency: Sonic and Hallo achieve strong lip-sync accuracy and emotional consistency, challenging unimodal detection. (3) Scalability: both frameworks support batch processing and domain-specific fine-tuning (e.g., celebrity voice dubbing), enabling rapid dataset expansion.
By integrating these methods, DigiFakeAV provides rich fake video and fake audio samples that closely resemble real-world forgery scenarios, thus supporting the development of robust deepfake detection systems.
3.3. Dataset Organization
Forged videos and audio can form four combinations: (1) Real Video–Real Audio (RV-RA), (2) Real Video–Fake Audio (RV-FA), (3) Fake Video–Real Audio (FV-RA), and (4) Fake Video–Fake Audio (FV-FA). We exclude RV-FA for two reasons: first, we focus on video-based facial forgery detection, whereas RV-FA does not involve human facial manipulation. For example, DFDC (dolhansky2020deepfakedetectionchallengedfdc, ) labels RV-FA samples as fake, which is inherently unfair to vision-only detectors. Second, audio-only manipulations without visual changes are often obvious and pose limited challenges for detection. For instance, FakeAVCeleb (khalid2022fakeavcelebnovelaudiovideomultimodal, ) has audio-video synchronization delays, which are relatively easy to detect.
We then introduce three combinations and the corresponding methods adopted, with detailed explanations:
Real Video–Real Audio (RV-RA). From nearly 40,000 clean video clips, we select 10,000 real videos representing diverse ethnicities, genders, and ages. These videos are further processed to crop portrait regions, forming the basis of our real video dataset.
Fake Video–Real Audio (FV-RA) This category contains fake videos generated by real audio-driven diffusion models. We first extract and convert audio from real videos into WAV format. RetinaFace (deng2019retinafacesinglestagedenseface, ) is then used to detect and crop representative facial frames. Using five digital human generation methods: Sonic (ji2024sonicshiftingfocusglobal, ), Hallo1 (xu2024hallohierarchicalaudiodrivenvisual, ), Hallo2 (cui2024hallo2longdurationhighresolutionaudiodriven, ), Echomimic (chen2024echomimiclifelikeaudiodrivenportrait, ), and V-Express (wang2024vexpressconditionaldropoutprogressive, ), we generate 25,000 forged videos.
Fake Video–Fake Audio (FV-FA). This category encompasses both forged audio and video. We employ a SOTA voice cloning method, CosyVoice 2 (du2024cosyvoice2scalablestreaming, ). Specifically, we first use a large language model to generate processed text. By providing real audio-text pairs, the CosyVoice 2 model synthesizes fake audio mimicking the target speaker’s voice characteristics. We then generate 25,000 forged videos based on this fake audio using algorithms like Sonic, Hallo1, Hallo2, and Echomimic.
3.4. Dataset Statistics
Previous datasets such as HDTF (zhang2021flow, ) and CelebV-HQ (zhu2022celebvhqlargescalevideofacial, ) exhibit demographic imbalances in data distributions. For example, HDTF contains a male-to-female ratio of 65:35, while CelebV-HQ includes predominantly Caucasian subjects. To address this, we show in Fig. 2 how we prioritize diversity across four dimensions during video selection: (1) Gender balance: we achieve a 57:43 male-to-female ratio to reflect real-world demographics and avoid excessive gender skew. (2) Ethnic diversity: the dataset includes White, African American, and Asian individuals, ensuring broad racial representation and enhancing model generalization. (3) Nationality and cultural background: subjects from diverse countries and cultural contexts are selected to reduce regional bias and enhance global representativeness. (4) Scenario diversity: video content simulates real-world scenarios, including social interactions, news broadcasts, and daily conversations, to improve applicability in practical scenarios.
3.5. User Study
To evaluate DigiFakeAV, we recruit 30 computer vision researchers, leveraging their domain expertise to assess forged videos. The study includes DigiFakeAV and three established datasets—FF++ (rössler2019faceforensicslearningdetectmanipulated, ), Celeb-DF (li2020celebdflargescalechallengingdataset, ), and DFDC (dolhansky2020deepfakedetectionchallengedfdc, ), which represent the first to third generations of deepfake video benchmarks.
We randomly select 50 video clips from each dataset and present them to participants via a dedicated evaluation interface. Participants are asked to judge each clip’s authenticity by voting whether it is real or fake.
As summarized in Tab. 3, the proposed DigiFakeAV dataset yields the highest misclassification rate among all tested datasets, with 68% of videos erroneously classified, comprising both ”real” (fake videos are classified as real) and ”fake” (real videos are classified as fake).
Table 3. Misclassification rates of videos labeled as ”fake” or ”real” in the user study across deepfake datasets.
Dataset FF++ Celeb-DF DFDC DigiFakeAV (ours) ”Fake” or ”real” (%) 12% 46% 32% 68%
4. DigiShield: A Multimodel Deepfake Detection Baseline
Existing detection models show substantial performance degradation on the DigiFakeAV dataset (detailed in the Sec. 5.2, Baseline Comparison), highlighting their limitations in capturing multimodal features of diffusion-based digital human forgeries. To address this, we introduce DigiShield, which enhances adaptability to novel forgery types and serves as a reproducible baseline and reference for future research.
4.1. Model Framework
Single-modality analysis proves inadequate for digital human forgery detection, making it essential to jointly model spatial and temporal inconsistencies. Building on prior spatiotemporal and two-stream networks (wang2023altfreezinggeneralvideoface, ; masi2020twobranchrecurrentnetworkisolating, ), we incorporate multimodal audio-visual cues. As shown in Fig. 4, our method includes (1) a spatiotemporal two-stream pipeline, (2) audio-visual fusion, and (3) a classification layer.
Spatiotemporal two-stream pipeline. Formally, given a set of video clips $x_{\raisebox{-0.90417pt}{\scriptsize in}}^{\raisebox{0.60275pt}{\scriptsize v}}\in\mathbb{R}^{T\times C\times H_{v}\times W_{v}}$ and audio signals $x_{\raisebox{-0.90417pt}{\scriptsize in}}^{\raisebox{0.60275pt}{\scriptsize a}}\in\mathbb{R}^{C\times H_{a}\times W_{a}}$ (where $T$ denotes the number of frames in the same video, $C$ represents the number of channels, $H_{v}$ $W_{v}$ indicates the spatial dimensions, The Mel-spectrogram $H_{a}\times W_{a}$ represents the frequency bins ($H_{a}$) and time frames ($W_{a}$) dimensions), we first employ a spatiotemporal two-stream pipeline based on 3D and 2D residual convolutions to capture forgery traces. The spatiotemporal representations of the video and audio can then be computed as follows:
(1) $f_{v}=\mathcal{F}_{v}\biggl{(}x_{\raisebox{-0.90417pt}{\scriptsize in}}^{\raisebox{0.60275pt}{\scriptsize v}}\biggr{)}\in\mathbb{R}^{T^{\prime}\times C^{\prime}\times H^{\prime}_{v}\times W^{\prime}_{v}},f_{a}=\mathcal{F}_{a}\biggl{(}x_{\raisebox{-0.90417pt}{\scriptsize in}}^{\raisebox{0.60275pt}{\scriptsize a}}\biggr{)}\in\mathbb{R}^{C^{\prime}\times H^{\prime}_{a}\times W^{\prime}_{a}}$
where $\mathcal{F}_{v}$ and $\mathcal{F}_{a}$ denote the visual and audio encoders respectively, the variables $T^{\prime}$, $C^{\prime}$, $H^{\prime}_{v}\times W^{\prime}_{v}$, and $H^{\prime}_{a}\times W^{\prime}_{a}$ denote the temporal, channel, spatial dimensions, and frequency-time representation of the video and audio features after convolutional and pooling operations, respectively.
Multimodal spatiotemporal fusion. The classification layer of visual and audio encoders is redesigned to produce spatiotemporal features for cross-modal attention fusion. Specifically, the extracted feature maps from both the video and audio streams, denoted as $f_{v}$ and $f_{a}$, are reshaped into flattened representations: $\widetilde{f}_{v}\in\mathbb{R}^{N_{v}\times d}$ and $\widetilde{f}_{a}\in\mathbb{R}^{N_{a}\times d},$ where $N_{v}=T^{\prime}\times H^{\prime}_{v}\times W^{\prime}_{v}$ and $N_{a}=H^{\prime}_{a}\times W^{\prime}_{a}$, with $d=C^{\prime}$ indicating the dimension of the feature channel.
We then employ cross-attention and self-attention modules to enhance multimodal interaction and feature fusion as follows:
(2) $\Psi$ $=\operatorname{CrossAtt}(\widetilde{f}_{v},\widetilde{f}_{a})=\operatorname{softmax}\left(\frac{\widetilde{f}_{v}W_{Q}(\widetilde{f}_{a}W_{K})^{\top}}{\sqrt{d_{k}}}\right)\widetilde{f}_{a}W_{V},$
$\Phi$ $=\operatorname{SelfAtt}(\Psi)=\operatorname{softmax}\left(\frac{\Psi W_{Q}^{\prime}(\Psi W_{K}^{\prime})^{\top}}{\sqrt{d_{k}}}\right)\Psi W_{V}^{\prime}.$
where $W_{Q},W_{K},W_{V}\in\mathbb{R}^{d\times d_{k}}$ are the learned projection matrices for queries, keys, and values in the cross-attention module; $W_{Q}^{\prime},W_{K}^{\prime},W_{V}^{\prime}\in\mathbb{R}^{d\times d_{k}}$ are the corresponding projection matrices in the self-attention module; $d_{k}$ is the dimension of the key (and query) vectors used for scaling the dot-product attention.
Final decision layer. After extracting audio and video features, the fused representation is formed by concatenating these multimodal features along the channel dimension, effectively integrating complementary information across modalities. These combined features are then processed through a series of downsampling (Down) modules to progressively reduce spatial redundancy while preserving salient characteristics. Finally, the refined representation is fed into a fully connected (FC) layer, which maps the feature embedding into a binary classification probability space, facilitating the final decision on the authenticity of the input video.
Figure 4. Overview of the DigiShield, it detects forgery traces by integrating spatiotemporal information from both video and audio modalities.
[…]
Overview of the DigiShield
4.2. Loss Function
The model is trained using both contrastive and cross-entropy losses to capture audio-visual inconsistencies indicative of forgery. To effectively fuse visual and audio modal features, we adopt a contrastive loss defined as:
(3) $\mathcal{L}_{con}=\frac{1}{N}\sum_{i=1}^{N}\big{[}$ $\ y_{i}\,D(\widetilde{\mathbf{h}}_{v}^{(i)},\widetilde{\mathbf{h}}_{a}^{(i)})^{2}$
$+(1-y_{i})\,\max\big{(}0,m-D(\widetilde{\mathbf{h}}_{v}^{(i)},\widetilde{\mathbf{h}}_{a}^{(i)})\big{)}^{2}\big{]}$
where $N$ is the batch size, $\widetilde{\mathbf{h}}_{v}^{(i)}$ and $\widetilde{\mathbf{h}}_{a}^{(i)}$ are the flattened visual and audio features of the $i$-th sample, $y_{i}\in\{0,1\}$ denotes ground-truth labels (1 for real, 0 for fake), $D(\cdot,\cdot)$ is the Euclidean distance, e.g., $D(\mathbf{x},\mathbf{y})=\|\mathbf{x}-\mathbf{y}\|_{2}$, and $m>0$ is a margin parameter.
The classification is guided by cross-entropy loss:
(4) $\mathcal{L}_{ce}=-\frac{1}{N}\sum_{i=1}^{N}\left[y_{i}\log\hat{y}_{i}+(1-y_{i})\log(1-\hat{y}_{i})\right]$
where $\hat{y}_{i}\in[0,1]$ is the predicted probability of the $i$-th sample being genuine.
The overall loss is a sum:
(5) $\mathcal{L}=\mathcal{L}_{con}+\mathcal{L}_{ce}$
This composite loss serves as the training objective.
5. Experiment
5.1. Implementation Details
During preprocessing, videos are processed with RetinaFace (deng2019retinafacesinglestagedenseface, ) for facial detection and cropping, while audio tracks are extracted, resampled to 16 kHz, and saved as WAV files. Mel-frequency cepstral coefficients (MFCCs) are computed for each frame and transformed into three-channel images as audio inputs for the multimodal model.
For DigiShield, we use 3D-ResNet-50 as the backbone. On DigiFakeAV [^4] Details of training setting are in the Appendix C, experiments adopt an 8:1:1 train/val/test split with no identity overlap, sampling 30 frames per video and applying color jitter and random cropping for augmentation. Heuristic upsampling is used to balance real and fake samples. The Area Under the Curve (AUC) is the primary evaluation metric, with all preprocessing and training settings kept consistent for fair comparison. We also use DF-TIMIT (korshunov2018deepfakesnewthreatface, ) and process the dataset following the official settings.
Table 4. AUC scores (%) of various methods on previous datasets and our DigiFakeAV.
Method DF-TIMIT FF-DF DFDC Celeb-DF FakeAVCeleb DigiFakeAV LQ HQ (ours) Meso4 87.8 68.4 84.7 75.3 54.8 60.9 50.1 MesoInception4 80.4 62.7 83.0 73.2 53.6 61.7 63.8 Xception-c23 95.9 94.4 99.7 72.2 65.3 72.5 66.1 Capsule 78.4 74.4 96.6 53.3 57.5 70.9 65.3 HeadPose 55.1 53.2 47.3 55.9 54.6 49.0 51.7 F3-Net 99.8 99.4 93.7 95.1 86.7 91.3 66.4 Cross Efficient ViT 50.4 55.8 99.1 95.1 86.7 80.5 52.2 SSVF - - - - - 94.5 51.0 SFIConv 100.0 100.0 95.9 96.7 95.8 93.0 71.2
5.2. Baseline Comparison
To conduct a thorough evaluation of the DigiFakeAV dataset, we systematically select nine representative deepfake detection algorithms based on the criteria of representativeness and accessibility. The selected models cover six categories: Naive (Meso4 (Afchar_2018, ), MesoInception4 (Afchar_2018, ), Xception-c23 (rössler2019faceforensicslearningdetectmanipulated, )), Spatial (Capsule (nguyen2019usecapsulenetworkdetect, ), HeadPose (yang2018exposingdeepfakesusing, )), Frequency (F3-Net (qian2020thinkingfrequencyfaceforgery, )), Temporal (Cross Efficient ViT (Coccomini_2022, )), Hybrid Domain (SFIConv (10286083, )), and Multi-modal (SSVF (feng2023selfsupervisedvideoforensicsaudiovisual, )).
As shown in Tab. 4, existing deepfake detection methods universally face performance collapse on the DigiFakeAV dataset. For example, the AUC of SFIConv drops from 100.0% (DF-TIMIT HQ) to 71.2%, and SSVF declines from 94.5% (FakeAVCeleb) to 51.0%. These results validate the disruptive challenge posed by diffusion-based digital human forgeries to existing detection paradigms. Traditional methods relying on local texture anomalies or static frequency-domain features are insufficient to handle highly coherent multimodal synthetic content, urgently requiring a reconstruction of the detection of spatiotemporal dynamics and cross-modal interactions [^5] Further evaluations of baseline comparison are provided in the Appendix D.
5.3. DigiShield Comparison
As shown in Tab.  5, DigiShield achieves an AUC of 80.1% on the highly challenging DigiFakeAV dataset, representing a 57.1% and 12.5% improvement over the current best multimodal detector SSVF (51.0%) and frequency-domain method SFIConv (71.2%), respectively. This performance breakthrough stems from the model’s targeted modeling of the intrinsic features of diffusion-based digital human forgeries. Meanwhile, DigiShield maintains a 100% AUC on the traditional face-swapping dataset DF-TIMIT, validating its generalization. Further analysis indicates that DigiShield’s success confirms the effectiveness of multimodal spatiotemporal inconsistency as a core metric for next-generation detection.
Table 5. Comparison of AUC scores for DigiShield on DigiFakeAV and DF-TIMIT
Method DigiFakeAV DF-TIMIT-LQ DF-TIMIT-HQ
MesoInception4 63.8 80.4 62.7
Capsule 65.3 78.4 74.4
Xception-c23 66.1 95.9 94.4
F3-Net 66.4 99.8 99.4
SFIConv 71.2 100.0 100.0
DigiShield (ours) 80.1 100.0 100.0
Table 6. Ablation study of DigiShield in DigiFakeAV.
$\mathcal{L}_{ce}$ $\mathcal{L}_{con}$ CrossAtt SelfAtt AUC
73.6
77.4
80.1
5.4. Ablation Study
Tab. 6 validates the effectiveness of our audio-visual spatiotemporal fusion strategy, we progressively introduce each component to assess its contribution. The baseline model uses a single-stream setting—by removing the audio extractor and training only with visual features under cross-entropy loss ($L_{ce}$)—which yields an AUC of 73.6%. This indicates that relying solely on single-modality classification is insufficient to capture inconsistency features. When the audio stream and contrastive loss ($L_{con}$) are incorporated to form a dual-stream architecture, and cross-modal attention (CrossAtt) is further introduced to explicitly model inter-modal discrepancies such as speech-lip asynchrony, the AUC is boosted to 77.4%. Finally, adding self-attention (SelfAtt) enhances long-term temporal coherence, pushing AUC to 80.1%. These results underscore the importance of both temporal coherence and cross-modal alignment in detecting digital human forgeries [^6] More details of the ablation study are in the Supplementary Materials (Appendix E).
6. Conclusion
DigiFakeAV is a new multimodal dataset based on advanced digital human forgery technologies. As a pioneering dataset emphasizing diffusion-based synthesis, it introduces unprecedented challenges for deepfake detection. We evaluate its difficulty using nine detection methods, showing significant performance degradation compared to previous benchmarks. Additionally, DigiShield, an audio-visual fusion baseline, is proposed to attempt to address this challenge. DigiFakeAV and DigiShield are expected to foster the development of more robust deepfake detectors and inspire new directions in detection research.